# STEM-PoM: Evaluating Language Models Math-Symbol Reasoning in Document Parsing

### v1

@@ 3.1 Setup
Table 4 shows second-level classification accuracy with full manuscript input. In this experiment, we assume that the model got the first-level classification correct. From the results, LSTM performs poorly, with an accuracy as low as 11.3% for predicting the DS. Larger models, like Llama2-13B and Mistral-8x7B, improve performance, especially in classifying Constants (up to 37.8%). Llama3-80B shows moderate improvements, with 40.2% accuracy for Global Operators, indicating reasonable capabilities in operator classification tasks. Claude3.5-Sonnet and GPT-3.5 show further improvements, particularly in Global Constants and Operators classification. GPT-3.5 achieves 48.5% accuracy for Local Constants and 49.7% for Global Operators. Lastly, GPT-4o provides the highest accuracy overall, reaching 60.5% for Local Operators and 58.6% for Matrix classification. By horizontally comparing the same model performance on different sub-attribute classifications, we find that the attribute Constants are generally easier to classify compared to Variables and Operators across all sizes of models, as seen by the overall higher accuracy in Constant-related tasks. However, Matrix and DS classification continue to present challenges, even for the largest models, indicating that certain structures or content types within manuscripts remain difficult to categorize accurately at the sub-attribute level.
@@
Overall, performance across all models on both first-level and second-level classification tasks shows a clear trend of improvement with increasing context length, highlighting the importance of context for accurately classifying mathematical symbols. Additionally, both small and large-size language models show a relatively higher accuracy in identifying Unit Descriptors and Operators compared to Variables and Constants, indicating that symbols with more distinct contextual or syntactical patterns are easier for models to classify. Through the above results, we aim to gain insights into the extent to which different category attributes of mathematical symbols influence LLMs’ understanding of math-rich documents by correctly classifying the symbols in real-world scenarios. We leave additional experiments in Appendix D.
@@ 4 Conclusion
In this paper, we introduce STEM-PoM, a comprehensive benchmark for evaluating language models’ mathematical reasoning abilities to classify math symbols from scientific texts. The dataset includes over 2,000 math instances sourced from ArXiv papers. Extensive experiments show that the best-performing model, achieves only 73.8% and 60.5% for first and second-level Part-of-Math Tagging classification accuracy, highlighting the challenge of extracting and categorizing mathematical symbols from large text corpora.
@@ Appendix A STEM-PoM Dataset Supplementary Materials
A.1 Frequency Analysis on math symbols
Figure 3: Total frequency (show-up times) of the top-50 mathematical symbols in the STEM-PoM. This illustrates the contextual polymorphism of a single mathematical symbol, i.e. it belongs to multiple different attribute categories depending on the related context or mathematical expression.
A.2 Source Dataset
MTDE [14] contains approximately 10,000 entries of mathematical symbol names along with their defined contexts. Each entry includes a ’short’ definition and a ’long’ definition. A short definition is a single-word definition, while a long definition consists of one or more words. The data was collected through random sampling from mathematical and scientific arXiv preprint manuscripts, covering a broad range of disciplines such as Physics, Computer Science, and Biology. For pre-processing, we ensured that the candidate data was generated via a corpus crawler and subsequently pruned and cleaned manually.
ArXiv Paper Dataset [22] contains 1.7 million arXiv articles, spanning a wide range of disciplines, including Mathematics, Physics, Chemistry, Economics, and Computer Science. We randomly sample 10,000 articles from this raw dataset and manually ensure that each manuscript is math-rich, containing numerous mathematical expressions and symbols. For pre-processing, we utilize the STEM-PoM Labeler to extract these symbols along with their surrounding context, ensuring that the data is representative of real-world mathematical usage across various scientific fields. Additionally, the extracted symbols and contexts are systematically cleaned and structured to facilitate further classification and analysis.
A.3 Dataset Definitions in Table 2
@@
A.4 First-Level and Second-Level Attributes Definition
Constant: A value that does not change in a mathematical expression.
Local Constant: Constant that is specific to a particular system or model, such as the gravitational constant in a simulation of a specific planetary system.
Global Constant: Constant that is applicable in all contexts, like the speed of light in a vacuum.
Discipline-specified Constant: Constant that applies to particular fields of study, for instance, Planck’s constant in quantum mechanics.
Operator: A symbol that operates on one or more operands.
Local Operators: Operator that is applied in a localized or specific context within a discipline, like a self-defined operation in matrix processing.
Global Operators: Operators that is used broadly across different disciplines, like the addition or multiplication operator.
Discipline-specified Operators: Operator that is unique to certain fields, such as the Hamiltonian operator in quantum physics.
Variable: A symbol that represents an unknown or changeable quantity in a mathematical expression.
Scalar: A quantity that has only magnitude, no direction.
Vector: A quantity that has both magnitude and direction.
Matrix: A rectangular array of numbers or symbols arranged in rows and columns.
A.5 STEM-PoM Labeler
@@
File Reading: We initiate the data importing operation progress by importing files from the designated arXiv folder, ensuring a structured and accessible repository of mathematical documents for subsequent processing.
Symbol Identification and Contextualization: For each file, we enumerate and display essential information: the current file being processed, the total number of symbols within, the sequence number of the current symbol, the graphical representation of the symbol, and the contextual content surrounding the symbol. This feature aids in providing a comprehensive overview and facilitates accurate symbol annotation.
Annotation Interface: We then present a user-friendly interface offering a set of predefined tagging options for each symbol. Through the designed interface, we easily select the most appropriate tag from these options, standardizing the labeling process and enhancing the consistency of the dataset.
Data Recording: Upon the selection of a tag for a symbol, We record this association by appending a new line to the dataset, capturing the symbol along with its assigned tag. This systematic data recording ensures the integrity and scalability of the MTCE dataset.
Dataset Evaluation: After constructing the dataset, we manually evaluate the quality and applicability of the annotated data. Specifically, we process the evaluation process through the following steps: Consistency Check, Inter-annotator Agreement, Statistical Analysis, and Benchmark Testing.
@@
Appendix B Backgrounds
Part-of-Math (PoM) Tagging : The part-of-math tagging task draws inspiration from similar tagging tasks such as part-of-speech tagging [36]. In the PoM context, the goal is to label individual mathematical tokens or expressions in math formulas with their corresponding mathematical roles. Such a task is essential for enabling a deeper semantic understanding of mathematical content by machines. Several datasets or benchmarks have been developed for the part-of-tagging task, but several challenges are still remaining. Abdou [43] collects mathematical content, such as formula representation and tagging for specific mathematical formula translations and verifications, including converting formulae into semantic LaTeX or testing with tools like CAS (Computer Algebra Systems). However, this focus on structured and narrow formula translations does not align with the broader, more diverse text-based tasks required to assess NLP models, due to the lack of scalability features in the collected math symbols. Ruocheng [37, 38] recently evaluated the potential of leveraging LLMs for automated annotation and Part-of-Math tagging of math symbols. However, their PoM tagging was conducted on the Digital Library of Mathematical Functions (DLMF) [27]. Since the source of math symbols is only one manuscript, the mathematical tokens collected only have a single classification type and are self-consistent. In contrast, our dataset incorporates the inherent messiness of published literature across several STEM subjects, where these domain-specific math symbols can have multiple classifications or meanings depending on the discipline and related context information.
Large Language Models: Pre-trained large language models (LLMs) have become a cornerstone in modern NLP [34, 46]. These models, which assign probabilities to word sequences by decomposing the probability of a sequence into the product of conditional probabilities of subsequent tokens, have evolved significantly over time. Early approaches were based on N-gram models, but with the advent of distributed word embeddings [5, 30], neural language models gained prominence. The scalability and performance improvements introduced by these models, along with the availability of vast textual data, have enabled the unsupervised pre-training of LLMs. These models, often referred to as foundation models [33, 23], can then be fine-tuned on smaller, task-specific datasets to adapt them for various downstream applications. For STEM-PoM, we apply one traditional sequence-based NLP model, LSTM [11], and several recent LLMs for our dataset evaluation.
Appendix C Additional Experiment Setups
@@
Training Details In our experiments, we train an LSTM with varying numbers of layers for the mathematical symbol classification tasks. For LLMs, we choose GPT-3.5 and apply a common parameter-efficient fine-tuning (PEFT) method, LoRA [16], to evaluate the model precision performance. We split our dataset into 80%/10%/10% for training/validation/testing sets.
@@
Model Parameters For the LSTM model, we use different layer sizes from {128, 256, 512, 1024}. The hidden state size is set to 256, the learning rate is set from {0.1, 0.01, 0.001}, the training epoch is 5, and the batch size is 16. We utilize the Adam optimizer [21]. For GPT-3.5 fine-tuning, we use the GPT-3.5-turbo-0125 model version and set the training epoch to 3. For LoRA fine-tuning, we set the LoRA rank to 32, batch size to 32, weight decay to 0.01, dropout to 0.1, and learning rate to $1e^{-4}$.
Appendix D Additional Experiments
D.1 Fine-tuning on STEM-PoM
@@
Table 5 shows the comparison result on main attributes between fine-tuned and directly vanilla-referenced GPT3.5. Notably, the fine-tuned GPT-3.5 model achieves an accuracy of 67.4% in the one-sentence context. However, its performance diminishes as the context length increases, with a noticeable drop to 66.9% for ten sentences and further down to 62.2% for full manuscript-length context. The decreasing trend suggests that while the fine-tuning process improves performance for shorter contexts, the model’s ability to handle longer contexts is hindered.
The vanilla inference results also show a similar pattern of improvement with context length, but the gap between fine-tuned and vanilla inference narrows as the context length grows. For instance, the difference in overall accuracy between fine-tuned and vanilla models is 10.6% for one-sentence contexts but only 1.6% for full manuscripts.
@@
Overall, the diminishing return for fine-tuned models with longer contexts indicates that fine-tuning amplifies sensitivity to the introduction of noisy or less relevant information when longer contexts are involved. The observation also could point to challenges in the fine-tuning process for long-context LLMs, which require more refined techniques to handle context length effectively.
D.2 Ablation Study
Model size(layers) Variable Constant Operator Unit Descriptor
128 24.5% 13.2% 10.3% 27.1%
256 28.7% 17.9% 15.7% 32.5%
512 34.2% 23.2% 24.9% 40.0%
1024 46.5% 35.9% 44.2% 51.3%
Context Length Variable Constant Operator Unit Descriptor
One Sentence 24.5% 13.2% 10.3% 27.1%
Five Sentence 26.3% 15.6% 14.1% 29.2%
Ten Sentence 28.1% 16.8% 15.5% 30.2%
Model Performance vs Model Size
Table 6 presents the classification accuracy of an LSTM model for first-level classification across different model sizes, ranging from 128 to 1024 layers. Note that we set the input context length to be one sentence. The results show a clear positive correlation between the model size and classification accuracy across all four categories. For the smallest model (128 layers), the accuracy ranges from 10.3% for the Operator class to 27.1% for the Unit Descriptor class. As the model size increases, the performance improves notably, with the largest model (1024 layers) achieving a relatively high-performance gain in accuracy, ranging from 35.9% for the Constant class to 51.3% for the Unit Descriptor class. The most substantial improvements are observed in the Operator category, where accuracy increases from 10.3% for 128 layers to 44.2% for 1024 layers. These results suggest that larger model sizes are more effective in capturing complex patterns.
Model Performance vs Data Input Lengths

### v2

@@ 4.2 First-Level Classification Results
Table 4 shows second-level classification accuracy with full manuscript input. In this experiment setting, we assume that the model got the first-level classification correct. By horizontally comparing the same model performance on different sub-attribute classifications, we find that the attribute Constants are generally easier to classify compared to Variables and Operators across all sizes of models, as seen by the overall higher accuracy in Constant-related tasks. However, Matrix and DS classification continue to present challenges, even for the largest models, indicating that certain structures, as well as content types within manuscripts, remain difficult to categorize accurately at the sub-attribute level.
@@
Overall, performance across all models on both first-level and second-level classification tasks shows a clear trend of improvement with increasing context length, highlighting the importance of context for accurately classifying mathematical symbols. Additionally, both small and large-size language models show a relatively higher accuracy in identifying Unit Descriptors and Operators compared to Variables and Constants, indicating that symbols with more distinct contextual or syntactical patterns are easier for models to classify. Through the above results, we aim to gain insights into the extent to which different category attributes of mathematical symbols influence LLMs’ understanding of math-rich documents by correctly classifying the symbols in real-world scenarios.
4.4 Fine-tuning on STEM-PoM
We fine-tune four different sizes of LLMs on STEM-PoM, as detailed in Table 5. Comparing these results with those in Table 3, we observe that additional training improves model classification accuracy for mathematical symbols when provided with more context information. However, continuously adding contextual knowledge to the training samples does not always lead to performance gains. We analyze this relationship further in Section 4.6.
4.5 Downstream Math Reasoning
After evaluating the performance of STEM-PoM across different LLMs, we further investigate its effectiveness in enhancing LLMs’ math reasoning capabilities. Specifically, we aim to address the following questions:
  • Q1: What is the relationship between an LLM’s ability to classify mathematical tokens and its real-world mathematical reasoning skills?
  • Q2: Does improving an LLM’s mathematical token classification capability enhance corresponding math reasoning abilities?
Models GSM8K MATH OlympiadBench Avg.
Llama2-13B 42.5% 29.1% 11.5% 27.7%
+ LoRA (STEM-PoM) 44.6% ($\uparrow 2.1$) 31.3% ($\uparrow 2.2$) 13.4% ($\uparrow 1.9$) 29.8% ($\uparrow 2.1$)
Mixtral-8x7B 72.4% 32.6% 13.7% 39.6%
+ LoRA (STEM-PoM) 74.1% ($\uparrow 1.7$) 34.1% ($\uparrow 1.5$) 16.4% ($\uparrow 2.7$) 41.5% ($\uparrow 1.9$)
Llama3.1-70B 91.6% 47.1% 26.4% 55.0%
+ LoRA (STEM-PoM) 93.2% ($\uparrow 1.6$) 48.8% ($\uparrow 1.7$) 28.2% ($\uparrow 1.8$) 56.7% ($\uparrow 1.7$)
GPT-4o 94.3% 88.7% 39.6% 74.2%
+ LoRA (STEM-PoM) 95.2% ($\uparrow 0.9$) 88.9% ($\uparrow 0.2$) 41.2% ($\uparrow 1.6$) 75.1% ($\uparrow 0.9$)
Table 6: Evaluation of LLMs on downstream mathematical reasoning tasks. + LoRA (STEM-PoM) indicates that the corresponding model is first fine-tuned using LoRA on STEM-PoM before being evaluated on downstream mathematical reasoning tasks. The improvements from LoRA fine-tuning on STEM-PoM are highlighted in darkgreen ($\uparrow x$).
To investigate these questions, we select challenging mathematical problems from GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), and OlympiadBench He et al. (2024) as our downstream tasks. We evaluate the model’s performance before and after LoRA fine-tuning on STEM-PoM. The evaluation is conducted using greedy decoding without tool integration. The primary metric reported is pass@1 accuracy with 3-shots CoT prompting. Full implementation details are shown in Appendix C.
Table 6 presents the experimental results on three mathematical reasoning datasets. A comparison across different models reveals that LLMs demonstrate superior performance on STEP-PoM and also generally achieve higher accuracy in mathematical reasoning tasks. This suggests that higher math token classification accuracy generally correlates with better mathematical reasoning performance. Additionally, all models exhibit performance gains after being fine-tuned on STEM-PoM. This suggests that enhancing LLMs’ understanding of mathematical symbols contributes to improved reasoning abilities in mathematical problem-solving. The experimental results align with the objective of our constructed dataset, which aims to fundamentally enhance LLMs’ mathematical reasoning through Part-of-Math Tagging.
4.6 Analysis on STEM-PoM
Model Performance vs Model Size.
Model size(layers) Variable Constant Operator Unit Descriptor
128 24.5% 13.2% 10.3% 27.1%
256 28.7% 17.9% 15.7% 32.5%
512 34.2% 23.2% 24.9% 40.0%
1024 46.5% 35.9% 44.2% 51.3%
Table 7 presents the classification accuracy of an LSTM model for first-level classification across different model sizes, ranging from 128 to 1024 layers. Note that we set the input context length to be one sentence. The results show a clear positive correlation between the model size and classification accuracy across all four categories. For the smallest model (128 layers), the accuracy ranges from 10.3% for the Operator class to 27.1% for the Unit Descriptor class. As the model size increases, the performance improves notably, with the largest model (1024 layers) achieving a relatively high-performance gain in accuracy, ranging from 35.9% for the Constant class to 51.3% for the Unit Descriptor class. The most substantial improvements are observed in the Operator category, where accuracy increases from 10.3% for 128 layers to 44.2% for 1024 layers. These results suggest that larger model sizes are more effective in capturing complex patterns. We further analyze the impact of input sequence length on model performance and compare the effectiveness of fine-tuning versus in-context learning on STEM-PoM. Please refer to D for full analysis.
Case Study and Further Discussion.
Please refer to Appendix E and F for the additional case study and discussions on STEM-PoM.
@@ 5 Conclusion
In this paper, we introduce STEM-PoM, a comprehensive benchmark for evaluating language models’ mathematical reasoning abilities to classify math symbols from scientific texts. The dataset includes over 2,000 math instances sourced from ArXiv papers. Extensive experiments show that the best-performing model, achieves only 73.8% and 60.5% for first and second-level classification accuracy, highlighting the challenge of extracting and categorizing math symbols from large text corpora.
6 Limitations
While we introduce a new benchmark dataset to evaluate LLMs’ reasoning abilities with math symbols, several challenges identified in our experimental evaluations warrant further investigation. First, as discussed in Section 4.3, even the most advanced LLMs struggle to accurately recognize complex math symbol attributes, such as matrices and discipline-specific constants or operators. To address this limitation, additional prompting or encoding methods, such as (Fu et al., 2023), need to be explored. Second, as evidenced by Table 3 and 4, when computational resources are constrained, leveraging smaller language models to process extensive mathematical context while maintaining performance levels comparable to larger models like GPT-4o presents a significant challenge. Through our benchmark, STEM-PoM, we intend to investigate these directions in future work to enhance LLMs’ mathematical reasoning capabilities within the scope of Part-of-Math Tagging.
7 Ethics Statement
The dataset proposed in this paper is derived from publicly available datasets that were released under licenses permitting reuse for research purposes. In constructing our dataset, we have taken care to ensure that all data sources comply with applicable legal and ethical standards, including privacy and data protection regulations. We have only used datasets that are free from personally identifiable information (PII) and sensitive attributes and have not modified or added any data that could compromise the anonymity or privacy of individuals. Furthermore, we encourage responsible use of the dataset and have made it available solely for academic and research purposes, with the expectation that future users will adhere to similar ethical guidelines.
@@ Appendix A STEM-PoM Dataset Supplementary Materials
A.1 Dataset Definitions in Table 2
@@
A.2 First-Level and Second-Level Attributes Definition
Constant: A value that does not change in a mathematical expression. Local Constant: Constant that is specific to a particular system or model, such as the gravitational constant in a simulation of a specific planetary system. Global Constant: Constant that is applicable in all contexts, like the speed of light in a vacuum. Discipline-specified Constant: Constant that applies to particular fields of study, for instance, Planck’s constant in quantum mechanics. Operator: A symbol that operates on one or more operands. Local Operators: Operator that is applied in a localized or specific context within a discipline, like a self-defined operation in matrix processing. Global Operators: Operators that is used broadly across different disciplines, like the addition or multiplication operator. Discipline-specified Operators: Operator that is unique to certain fields, such as the Hamiltonian operator in quantum physics. Variable: A symbol that represents an unknown or changeable quantity in a mathematical expression. Scalar: A quantity that has only magnitude, no direction. Vector: A quantity that has both magnitude and direction. Matrix: A rectangular array of numbers or symbols arranged in rows and columns.
A.3 STEM-PoM Labeler
@@
File Reading.
We initiate the data importing operation progress by importing files from the designated arXiv folder, ensuring a structured and accessible repository of mathematical documents for subsequent processing.
Symbol Identification and Contextualization.
For each file, we enumerate and display essential information: the current file being processed, the total number of symbols within, the sequence number of the current symbol, the graphical representation of the symbol, and the contextual content surrounding the symbol. This feature aids in providing a comprehensive overview and facilitates accurate symbol annotation.
Annotation Interface.
We then present a user-friendly interface offering a set of predefined tagging options for each symbol. Through the designed interface, we easily select the most appropriate tag from these options, standardizing the labeling process and enhancing the consistency of the dataset.
Data Recording.
Upon the selection of a tag for a symbol, We record this association by appending a new line to the dataset, capturing the symbol along with its assigned tag. This systematic data recording ensures the integrity and scalability of the MTCE dataset.
@@
Appendix B Additional Experiment Setups
@@
Training Details. In our experiments, we train an LSTM with varying numbers of layers for the mathematical symbol classification tasks. For LLMs fine-tuning, we apply the common parameter-efficient fine-tuning (PEFT) method, LoRA Hu et al. (2021), to evaluate the model precision performance. Specifically, we set the LoRA rank to 32, batch size to 32, weight decay to 0.01, dropout to 0.1, and learning rate to $1e^{-4}$.
Dataset Split.
We divide STEM-PoM into 80%/10%/10% for training/validation/testing sets.
Model Details.
@@
For the LSTM model, we use different layer sizes from {128, 256, 512, 1024}. The hidden state size is set to 256, the learning rate is set from {0.1, 0.01, 0.001}, the training epoch is 5, and the batch size is 16. We utilize the Adam optimizer Kingma (2014). For LLMs generation, we set the temperature for all models to 0 (Greedy Decoding), top_p to 1.0, frequency penalty to 0, and presence penalty to 0.
Appendix C Downstream Math Reasoning
Evaluation Details. We use a unified 3-shot CoT prompt template for all models, provided in Figure 7. Decoding is performed with a temperature of 0 (Greedy Decoding), and we report pass@1 accuracy. Dataset details are outlined below.
  • GSM8K Cobbe et al. (2021) is a mathematical dataset comprising 8.5K high-quality, linguistically diverse grade-school math word problems designed for multi-step reasoning. Solutions involve elementary arithmetic operations and require no concepts beyond early algebra. The test set consists of 1,319 unique problems.
  • MATH Greiner-Petter et al. (2022) is a dataset of 12,500 challenging competition-level mathematics problems sourced from contests such as AMC 10, AMC 12, and AIME. Each problem is accompanied by a step-by-step solution, enabling models to learn answer derivations and explanations. The test set comprises 5,000 unique problems.
  • OlympiadBench He et al. (2024) is a bilingual, multimodal scientific benchmark comprising 8,476 Olympiad-level math and physics problems, including those from the Chinese college entrance exam. For our evaluation, we use the open-ended, text-only math competition subset, which consists of 674 problems.
Implementation Details.
For LLMs generation, we set the temperature for all models to 0 (Greedy Decoding), top_p to 1.0, frequency penalty to 0, and presence penalty to 0.
Appendix D Full Analysis on STEM-PoM
Context Length Variable Constant Operator Unit Descriptor
One Sentence 24.5% 13.2% 10.3% 27.1%
Five Sentence 26.3% 15.6% 14.1% 29.2%
Ten Sentence 28.1% 16.8% 15.5% 30.2%
Model Performance vs Data Input Lengths.
@@ Appendix D Full Analysis on STEM-PoM
Fine-tuning vs In-context learning.
@@
Table 9 shows the comparison result on main attributes between fine-tuned and directly vanilla-referenced GPT3.5. Notably, the fine-tuned GPT-3.5 model achieves an accuracy of 67.4% in the one-sentence context. However, its performance diminishes as the context length increases, with a noticeable drop to 66.9% for ten sentences and further down to 62.2% for full manuscript-length context. Such diminishing return for fine-tuned models with longer contexts indicates that fine-tuning amplifies sensitivity to the introduction of noisy or less relevant information when longer contexts are involved. The observation also could point to challenges in the fine-tuning process for long-context LLMs, which require more refined techniques to handle context length effectively.
Appendix E Case Study on STEM-PoM
To further demonstrate the utility and effectiveness of the STEM-POM dataset, we conducted a case study focusing on error-prone classification scenarios involving mathematical symbols with ambiguous or context-dependent attributes. Our main goal for this case study is to evaluate where and why current state-of-the-art language models fail in complex mathematical reasoning tasks.
E.1 Setup
Sub-dataset Selection.
Building on the main experimental results, we select a subset of mathematical symbols from STEM-PoM dataset—particularly those frequently misclassified—to analyze the model’s failure modes.
Evaluation Metrics.
We focused on precision accuracy and the frequency of specific errors across the selected symbols. Additionally, we conducted qualitative error analysis to identify consistent failure cases.
Experimental Contexts.
Two context lengths were considered for testing:
  • Ten Sentences: Top Ten sentences selected by human experts surrounding the symbol.
  • Full Manuscript: Entire input manuscript.
E.2 Results and Analysis
Error Frequency and Common Failure Cases.
Table 10 highlights the top ten symbols that LLMs frequently misclassified. The error rates are calculated as the percentage of incorrect predictions across all contexts. Notably, even the best-performing model, GPT-4o, exhibited a significant error rate for symbols with overlapping roles, such as "$\alpha$" and "$\Sigma$".
Symbol GPT-4o Claude 3.5 Llama3.1-70B Mistral-8x7B
$\alpha$ 25% 32% 45% 58%
$\beta$ 22% 29% 38% 50%
$\Delta$ 30% 35% 48% 60%
$\psi$ 28% 33% 44% 57%
$\theta$ 26% 31% 42% 55%
$\Sigma$ 34% 38% 50% 63%
$\lambda$ 20% 27% 36% 48%
$\pi$ 18% 25% 34% 46%
$\mu$ 24% 30% 41% 54%
$\xi$ 27% 34% 46% 59%
Table 10: Top 10 error-prone symbols and their misclassification rates.
Error Analysis by Symbol Type.
The analysis revealed recurring error patterns: (1) Ambiguity in Context: Symbols like "$\alpha$" and "$\beta$" were frequently misclassified due to their diverse meanings in different disciplines. (2) Overlapping Attributes: "$\Sigma$" and "$\Delta$" were often confused between operators and constants depending on context. (3) Domain-Specific Knowledge Gaps: Models struggled with symbols like "$\psi$" and "$\theta$" that require understanding of physics or quantum mechanics.
Symbol Contextual Sentence Incorrect Prediction (Correct)
$\alpha$The coefficient $\alpha$ controls the rate of decay in the model." Constant (Correct: Variable)
$\psi$The wave function $\psi$ represents the quantum state of the system." Variable (Correct: Operator)
$\Sigma$The summation $\Sigma$ is over all possible states." Operator (Correct: Constant)
Table 11: Qualitative analysis of errors for selected symbols.
Performance Comparison.
Figure 5 illustrates the misclassification rates for the top ten error-prone symbols across models. GPT-4o consistently outperformed smaller models, but the error rates remained substantial for challenging symbols.
Qualitative Examples.
Table 11 provides examples of contextual sentences and the corresponding model predictions. These examples demonstrate how insufficient understanding of context leads to systematic errors. Symbols like "$\alpha$" and "$\delta$" are often misclassified due to ambiguous contextual cues. For example, "$\alpha$" is misclassified as a constant in differential equations, where it actually serves as a prefix operator given the context.
Impact of Context Length.
While longer contexts improve classification accuracy overall, they also introduce noise. For example, in the full manuscript setting, models sometimes overfit on irrelevant context, leading to errors in identifying local operators as global constants.
Figure 5: Misclassification rates for the top 10 error-prone symbols across different LLMs.
Through the case study, we conclude that despite the contextual advantages provided by STEM-POM, symbols with ambiguous or multifaceted roles remain challenging. Larger models like GPT-4o mitigate some of these issues but still fail to handle nuanced sub-attributes effectively.
Appendix F Discussions
Impact and Applicability of STEM-PoM on Math Reasoning.
In this section, we discuss why classifying mathematical tokens is highly relevant to advancing mathematical reasoning on LLMs.
On the surface level, summarization is a quintessential AI task, and being able to summarize mathematical tokens is a crucial downstream task that has already been held back in several recent research. One notable example is ScholarPhi Head et al. (2021), where researchers endeavor to make STEM manuscripts more accessible by enhancing a PDF viewer with an annotation tool that summarizes mathematical tokens and equations whenever users click on them. The authors observed that high-quality annotations significantly enhanced users’ comprehension of the manuscripts. However, they manually annotated their test sets, as they were uncertain about how annotation accuracy would influence comprehension.
Furthermore, researchers have explored the task of converting mathematical content in manuscripts into computable functions Greiner-Petter et al. (2022); Greiner-Petter (2023). In such tasks, where mathematical expressions must be transformed into alternative formats, accurately classifying these tokens is crucial.
Therefore, the classification of mathematical tokens serves as a foundational component for numerous downstream tasks. We developed this dataset to provide researchers with a baseline for evaluating and advancing this critical functionality.
Appendix G Model Prompts
For detailed prompts of STEM-PoM and downstream math reasoning tasks, please refer to Figure 6 and 7.
Figure 6: The prompt design for STEM-PoM.
Figure 7: The prompt design for downstream math reasoning tasks including GSM8K, MATH, and OlypiadBench.